The Ball Corporation's beverage can manufacturing plant in Fort Atkinson, Wisconsin, uses a metal supplier that provides metal with a known thickness standard deviation σ = .000943 mm. Assume a random sample of 58 sheets of metal resulted in an x¯ = .2603 mm. Calculate the 99 percent confidence interval for the true mean metal thickness. (Round your answers to 4 decimal places.) The 99% confidence interval is from to
Answer #1

Formula for Confidence Interval for population Mean When Population Standard deviation is known

Population Standard deviation : : metal with a known thickness standard deviation = 0.000943 mm

Sample size : n: Number of sheets of metal in the sample = 58

Sample mean : = 0.2603

for 99% confidence level = (100-99)/100 = 0.01

/2 = 0.01/2 = 0.005

99 percent confidence interval for the true mean metal thickness

99% confidence interval is from 0.2600 to 0.2606